Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Saturday, November 7, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 7, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/07/98 (Saturday)

Segment 1 (Emidio to ANF) Report for 11/07/98. Trench was excavated at the Tejon Bench, moving southward to ~EMP 18.50. Immediately north of the excavator, a curved section of pipe was lowered into the trench for a tie-in, ( ~EMP 18.45). Shading of pipe and padding operations occurred between EMP 16.90 and 17.06. Trench was opened for a tie-in to be made at EMP 16.89. Welding, NDT and Coating crews worked on a 15- section length of pipe at EMP .30. A weld was completed and coated at the bore pit along Rock Road (~EMP 8.95). Electrical work and station construction continued at Emidio and Grapevine stations.

Segment 2B (Castaic to Sylmar): Report for 7 November 1998: The directional bore for the fiber optic conduit under the Santa Clara River kicked off today. The crew had a late start, and that coupled with equipment difficulties, resulted in slow progress.

Segment 3B & 3C (Sylmar to Los Angeles): Report for 7 November 1998: 3B: Crews worked on clean up of the ROW just north of the Downey Bridge in the MTA maintenance facility. 3C: The bore crew at Hubbard Avenue pushed 40 feet of casing pipe today. Crews worked at the California Aqueduct removing frame work from the concrete pour.

Segment 4B (El Segundo to Lynwood) 11-7-98: Backfilling activities continued at 108 Th St. ESEMP 1.11. The crew at Santa Ana and Alameda ESEMP 0.04. was paving today. Construction activities continued at El Segundo station ESEMP 11.00. This crew was pushing the pipe through the casing. The crew at Slater ESEMP 1.68 was repairing the conduit for the fiberoptics. The crew at Lynwood station EMP 0.0. was doing the hydro test inside the station.

Segment 5 (Lynwood to Long Beach) Report for 7 November 1998: Problems temporarily halted the reaming of the directional drill at Highway 91/Compton Creek. The casing bore was completed at Wilmington station and the crew is beginning to push the pipe through.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page